# Holstein-Friesian Re-Identification using Multiple Cameras and Self-Supervision on a Working Farm

### v2

@@ 2.1 Re-Identification and Farm Systems
Learning in Dynamic Data Domains. Adaptive transfer learning and self-supervised learning have aided address such dataset-wide scenarios by transferring meaningful knowledge from an existing model onto unlabelled datasets that retain some structure (e.g. the same animal species – even if patterned completely differently). Reliant on a ResNet backbone and spectral embedding analysis, Able to propose accurate identification over smaller batches of data efficiently, Pathak et al. [16] refined earlier human Re-ID tools through inner fine-graining across spatial subregions of latent space. Therien et al. [17] then extended this Re-ID framework for 3D depth sensor applications by utilising PointNet and ViT. We note that these approaches are, in principle, fully applicable to the Holstein-Friesian cattle Re-ID task.
@@ 2.2 Species Recognition
CNN-based Object Recognition. Deep residual CNNs and their derivatives are the most widespread tool-kits for semantic detection and segmentation. The introduction of region proposal networks (RPNs) allowed for performing such tasks faster and perform real-time segmentation tasks. Later, pyramid-shaped networks utilised stacking region selection with inter-layer knowledge to acquire multi-level features. Knowledge of the previously trained semantic segmentation model can also be further distilled for transfer learning over advanced and detailed features. Based on these and related approaches, animal recognition and even behaviour analysis has not only been applied to wild species [21], but also to Holstein-Friesian cattle [22] including work by Andrew et.al. [12] proposing a drone-based, automatic CNN framework to monitor Holstein-Friesian cows from the air.
@@ 3.2 Tracklet Generation and Refinement
Species Detection and Tracklet Generation. Adapted from previous works by Gao et.al.[14], we first applied species detection [13] to the farm footage in order to extract cattle locations for each frame in the CCTV video (see Fig. 3). We trained this species detector for cow torso recognition using pre-trained weights from ImageNet for ResNet-152 as our backbone. We adapted the tools from NVIDIA Computer Vision Toolkit [^1] https://github.com/NVIDIA/retinanet-examples to perform head-oriented, rotated bounding box detections, which are later normalised so that all images to have the same orientation – i.e. the cow is always seen walking left-to-right in the dataset. Due to the wide visual range of views, we limited detection to a region of interest (RoI) in the cattle lane for each camera (see Fig. 1). Finally, we fine-tuned our species detector via Adam optimiser using $5,606$ annotated frames across three selected views. Fig. 4 shows a training and performance overview for this species detector.
@@ 3.3 Cattle Re-Identification
Loss Function, Evaluation, and Experimental Setup. Within the supervised framework, we employ a model incorporating cross-entropy loss and a one-hot classifier on top of the reshaped images. To address variations in quantity amongst individuals (i.e. class imbalances), we sample images of individuals with equal probability for training. Within the self-supervised framework, a linear layer was utilised to generate multi-dimensional feature vectors for metric learning purposes. The NTXentLoss [26] was used as a contrastive loss function. It encourages distance reduction of items having the same label in latent space while moving these items away from those with contrasting labels. To ensure that the ground-truth identities of the cows could not ‘leak’ into the self-supervision process, pseudo-labels were used (effectively randomizing the order in which cows appeared daily). For testing our unsupervised classifier, the ground-truth labels were used and KNN clustering was applied to evaluate performance. As part of the comparative analysis, we also substituted the ImageNet weights from our frozen self-supervision counterpart into the supervised classifier to analyse the model’s ability to learn the intricate mapping between cow body feature maps and their corresponding labels. Both the supervised and self-supervised frameworks have been trained using NVIDIA GeForce RTX 2080 Ti GPUs.
@@ 4 Results
Re-Identification Performance. Here we present results on the task of individual identification based on training from tracklets as described above. Table 1 represents the quantitative performance of our supervised classifier and self-supervised identifier, respectively, where we compared the results of training for each independent camera and also with them all combined. For our classifier trained via supervised methods, the performance of merging the camera views reached the best accuracy result at $96.06\%$ with the cropped image size being $128\times 128$ using a $32$D-embedding. This is marginally inferior to camera $1$ performance ($96.78\%$), but is significantly improved compared to previously published single camera performance in slightly different settings benchmarked at $76.9\%$ accuracy as reported for Cows2021 [14]. Self-supervision demonstrated a small overall increase in performance when merging multiple views together as shown in Table 1. When training with our ResNet18 backbone with images resized to $128\times 128$, combining views leads to an accuracy of $96.36\%$, boosting the performance of camera $1$, $2$ and $3$ by $0.48\%$, $3.62\%$, and $39.67\%$ respectively. With more images causing dataset imbalance, the relatively large gap between the performance of camera $3$ versus the combined views was due to the cows in camera $3$ being mostly stationary (they tend to stop walking after the milk-race, so group together making it difficult to disambiguate individuals accurately). Cattle detection in dense groups (due to a camouflaging effect of the coat pattern) is a truly difficult task not solved to the same quality levels as standard object detection. Fig. 7 and A.2 provide visualisations of embeddings obtained during inference from self-supervision. Additionally, in Fig. 6 we show the confusion matrices of ground truth vs. prediction. Note that pattern similarity remains the major reason for misidentification. For instance, cow $035$ is often misclassified as cow $031$ due to the similarity of their body patterning from the rear and the side.
@@ 5 Discussions
6 Conclusions and Future Work
We introduced MultiCamCows2024, the first dataset consisting of multi-day images of Holstein-Friesian cows covered by multiple distinct cameras in a working farm. We performed supervised and self-supervised experiments using both merged and single cameras alone to test if fusing camera views enhances cow re-identification. We found that within our closed-set, fusing data across cameras boosts the performance of the cattle re-identification task over that of simply using each view independently (assuming the same level of data utilisation). We now discuss how our work may be further extended.
@@ 6 Conclusions and Future Work
Appendix A
A.1 Dataset Metadata
@@
A.2 Identifier Embedding and Metrics Visualisation

### v3

@@ 2.1 Re-Identification and Farm Systems
Learning in Dynamic Data Domains. Adaptive transfer learning and self-supervised learning have aided address such dataset-wide scenarios by transferring meaningful knowledge from an existing model onto unlabelled datasets that retain some structure (e.g. the same animal species – even if patterned completely differently). Reliant on a ResNet backbone and spectral embedding analysis and able to propose accurate identification over smaller batches of data efficiently, Pathak et al. [16] refined earlier human Re-ID tools through inner fine-graining across spatial subregions of latent space. Therien et al. [17] then extended this Re-ID framework for 3D depth sensor applications by utilising PointNet and ViT. We note that these approaches are, in principle, fully applicable to the Holstein-Friesian cattle Re-ID task.
@@ 2.2 Species Recognition
CNN-based Object Recognition. Deep residual CNNs and their derivatives are the most widespread toolkits for semantic detection and segmentation. The introduction of region proposal networks (RPNs) allowed for performing such tasks faster and perform real-time segmentation tasks. Later, pyramid-shaped networks utilised stacking region selection with inter-layer knowledge to acquire multi-level features. Knowledge of the previously trained semantic segmentation model can also be further distilled for transfer learning over advanced and detailed features. Based on these and related approaches, animal recognition and even behaviour analysis has not only been applied to wild species [21], but also to Holstein-Friesian cattle [22] including work by Andrew et.al. [12] proposing a drone-based, automatic CNN framework to monitor Holstein-Friesian cows from the air.
@@ 3.2 Tracklet Generation and Refinement
Species Detection and Tracklet Generation. Adapted from previous works by Gao et.al.[14], we first applied animal detection [13] to the farm footage in order to extract cattle locations for each frame in the CCTV video (see Fig. 3). We trained this species detector for cow torso recognition using pre-trained weights from ImageNet for ResNet-152 as our backbone. We adapted the tools from NVIDIA Computer Vision Toolkit [^1] https://github.com/NVIDIA/retinanet-examples to perform head-oriented, rotated bounding box detections, which are later normalised so that all images to have the same orientation – i.e. the cow is always seen walking left-to-right in the dataset. Due to the wide visual range of views, we limited detection to a region of interest (RoI) in the cattle lane for each camera (see Fig. 1). Finally, we fine-tuned our species detector via Adam optimiser using $5,606$ annotated frames across three selected views. Fig. 4 shows a training and performance overview for this species detector.
@@ 3.3 Cattle Re-Identification
Loss Function, Evaluation, and Experimental Setup. Within the supervised framework, we employ a model incorporating cross-entropy loss and a one-hot classifier on top of the reshaped images. To address variations in quantity amongst individuals (i.e. class imbalances), we sample images of individuals with equal probability for training. Within the self-supervised framework, a linear layer was utilised to generate multi-dimensional feature vectors for metric learning purposes. The NTXentLoss [26] was used as a contrastive loss function. It encourages distance reduction of items having the same label in latent space while moving these items away from those with contrasting labels. To ensure that the ground-truth identities of the cows could not ‘leak’ into the self-supervision process, pseudo-labels were used (effectively randomizing the order in which cows appeared daily). For testing our unsupervised classifier, the ground-truth labels were used and KNN clustering was applied to evaluate performance. Both the supervised and self-supervised frameworks have been trained using NVIDIA GeForce RTX 2080 Ti GPUs.
@@ 4 Results
Re-Identification Performance. Here we present results on the task of individual identification based on training from tracklets as described above. Table 1 represents the quantitative performance of our supervised classifier and self-supervised identifier, respectively, where we compared the results of training for each independent camera and also with them all combined. For our classifier trained via supervised methods, the performance of merging the camera views reached the best accuracy result at $96.06\%$ with the cropped image size being $128\times 128$ using a $32$D-embedding. This is marginally inferior to camera $1$ performance ($96.78\%$), but is significantly improved compared to previously published single camera performance in slightly different settings benchmarked at $76.9\%$ accuracy as reported for Cows2021 [14]. Self-supervision demonstrated a small overall increase in performance when merging multiple views together as shown in Table 1. When training with our ResNet18 backbone with images resized to $128\times 128$, combining views leads to an accuracy of $96.36\%$, boosting the performance of camera $1$, $2$ and $3$ by $0.48\%$, $3.62\%$, and $39.67\%$ respectively. With more images causing dataset imbalance, the relatively large gap between the performance of camera $3$ versus the combined views was due to the cows in camera $3$ being mostly stationary (they tend to stop walking after the milk-race, so group together making it difficult to disambiguate individuals accurately). Cattle detection in dense groups (due to a camouflaging effect of the coat pattern) is a truly difficult task not solved to the same quality levels as standard object detection. Fig. 7 and B provide visualisations of embeddings obtained during inference from self-supervision. Additionally, in Fig. 6 we show the confusion matrices of ground truth vs. prediction. Note that pattern similarity remains the major reason for misidentification. For instance, cow $035$ is often misclassified as cow $031$ due to the similarity of their body patterning from the rear and the side.
@@ 5 Discussions
5.1 Future Works
@@
6 Conclusion
We introduced MultiCamCows2024, the first dataset consisting of multi-day images of Holstein-Friesian cows covered by multiple distinct cameras in a working farm. We performed supervised and self-supervised experiments using both merged and single cameras alone to test if fusing camera views enhances cow re-identification. We found that within our closed-set, fusing data across cameras boosts the performance of the cattle re-identification task over that of simply using each view independently (assuming the same level of data utilisation).
Appendix A Dataset Metadata
@@
Appendix B Identifier Embedding and Metrics Visualisation